OpenStudy (anonymous):

Multiply: (3a + 11)(3a - 11) im having trouble with these stuff. i get it and everything i just got confused and i dont know if i got the right answer. i think the right answer is 9a^2 -121. Is it wrong?

Parth (parthkohli):

As I said in the last post - \((a + b)(a - b) = a^2 - b^2\) Here it'd be \((3a + 11)(3a - 11) = (3a)^2 - (11)^2\)

Parth (parthkohli):

http://openstudy.com/updates/4fd62be0e4b04bec7f17374c See this tutorial by me, this may help. :)

OpenStudy (anonymous):

\[9a^2 - 121\] is the right answer...

OpenStudy (lgbasallote):

if you doubt then go back to good old FOIL :) (3a x 3a) + (3a x -11) + (11 x 3a) + (11 x -11) 9a^2 - 33a + 33a - 121 9a^2 - 121! :D

OpenStudy (anonymous):

do FOIL! ....first outside, inside, last. 9a^2 - 33a + 33a - 121 then simplify it to 9a^ 2 - 121